# Rethinking the Value of Agent-Generated Tests for LLM-Based Software Engineering Agents

### v1

@@ Abstract.
Large Language Model (LLM) code agents increasingly resolve repository-level issues by iteratively editing code, invoking tools, and validating candidate patches. In these workflows, agents often write tests on the fly, a paradigm adopted by many high-ranking agents on the SWE-bench leaderboard. However, we observe that GPT-5.2, which writes almost no new tests, can even achieve performance comparable to top-ranking agents. This raises the critical question: whether such tests meaningfully improve issue resolution or merely mimic human testing practices while consuming a substantial interaction budget.
To reveal the impact of agent-written tests, we present an empirical study that analyzes agent trajectories across six state-of-the-art LLMs on SWE-bench Verified. Our results show that while test writing is commonly adopted, but resolved and unresolved tasks within the same model exhibit similar test-writing frequencies Furthermore, these tests typically serve as observational feedback channels, where agents prefer value-revealing print statements significantly more than formal assertion-based checks. Based on these insights, we perform a controlled experiment by revising the prompts of four agents to either increase or reduce test writing. The results suggest that changes in the volume of agent-written tests do not significantly change final outcomes. Taken together, our study reveals that current test-writing practices may provide marginal utility in autonomous software engineering tasks.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2026; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Code agents are increasingly used as effective paradigm for resolving software issues, where Large Language Models (LLMs) (OpenAI, 2025; DeepSeek, 2025b; Chen and Jiang, 2024) are integrated into scaffold of tools and interaction protocols to edit real repositories, invoke tools, and attempt to resolve issues end-to-end (Yang et al., 2024a; Örwall, 2024; Wang et al., 2024b; Gao et al., 2025b; Hong et al., 2024; Qian et al., 2024; Ehrlich et al., 2025; Xia et al., 2024). In this paper, a code agent denotes an LLM coupled with external tools and an iterative action–observation loop, and the scaffold refers to the surrounding tool interface and interaction protocol that specifies the agent’s allowed actions and feedback. Among the diverse skills required by code agents, testing plays a critical role, which expose regressions, validate hypotheses, and provide a feedback loop during patch development (Zhang et al., 2024; Xia et al., 2024; Liu et al., 2024; Mu et al., 2025).
Figure 1. Overview of our study design. RQ1 profiles emergent testing behaviours (test writing frequency, timing, and execution analysis). RQ2 characterizes the feedback signals encoded in agent-written tests (assertions vs. value-revealing prints) and the types of assertions. RQ3 applies prompt interventions to encourage or discourage writing tests, and measures both outcome impact and efficiency impact.
@@
To better understand agent-written tests, we conduct a preliminary quantitative analysis of agent trajectories on SWE-bench Verified (OpenAI, 2024) using mini-SWE-agent (SWE-agent Team, 2024), where testing is optional and not enforced by any hard-coded procedure. We find that agent-written testing is prevalent for several strong models. For example, claude-opus-4.5 (ranked #1 in this setting, 74.4% resolution) generates at least one new test artifact in about 83% of tasks. Surprisingly, we observe a pronounced contrast for gpt-5.2: it achieves a comparable resolution rate (71.8%), only 2.6 percentage points below claude-opus-4.5, while generating near-zero new tests (only in 0.6% of tasks). This interesting observation inspires us to a core question: Do agent-written tests truly facilitate task resolution, or do models merely mimic a learned software development practice while the resulting tests contribute little to the final patch? If the latter holds, then the widespread creation and execution of agent-written tests may represent a significant resource waste, consuming significant interaction budget without meaningful gains in task success. Therefore, we argue that a systematic empirical study is needed to understand the role of agent-written tests in resolving software issues.
Prior work mostly evaluates and benchmarks LLM-generated tests under predefined testing objectives and fixed quality metrics (e.g., unit tests, assertions, or issue-reproducing tests), typically with respect to a fixed target program or snapshot of code under test (Lops et al., 2025; Zhang et al., 2025b; Mündler et al., 2024; Yuan et al., 2024; Wang et al., 2025, 2024a; Schäfer et al., 2023). However, in complex real-world GitHub issue resolution (Jimenez et al., 2024; Khandpur et al., 2025), the codebase and candidate patches evolve over time, and test writing and usage arise dynamically as self-directed behaviors rather than pre-specified evaluation objectives. Yet the intrinsic tendency of high-autonomy agents to write and use tests during such issue resolution—and the impact of such agent-written tests on resolution outcomes—has not been systematically studied. This motivates a closer empirical investigation of agent-written tests, guided by the following three research questions.
@@
Guided by the above gap, we conduct a systematic empirical study of agent-written tests in GitHub issue resolution. Figure 1 summarizes our study design, which decomposes the problem into three complementary research questions. RQ1 characterizes the agents’ testing behaviors when test writing is required by the prompt: whether agents write tests, when they introduce them, and how intensively they execute them. RQ2 shifts from behaviors to test content, investigating what feedback signals agent-written tests actually emit at execution time (assertions vs. value-revealing prints) and what types of assertions agents use. RQ3 evaluates impact: by revising prompts to encourage or discourage writing new tests, we measure whether changing test-writing behavior meaningfully alters task resolution outcomes and what efficiency costs (API calls and tokens) these changes incur.
@@
Across models, agent-written testing is best understood as a model-dependent process style rather than a dependable driver of success. First, RQ1 shows that test writing is a widespread behavior among the studied models, while resolved and unresolved trajectories within a model exhibit broadly similar test-writing rates. When agent-written tests exist, unsuccessful trajectories tend to spread test writing slightly more over the run and execute tests more frequently. Second, RQ2 reveals that agent-written tests primarily function as an observational feedback channel: value-revealing prints consistently dominate assert-based checks, and assertion forms concentrate on local-property and exact-value checks, with relational/range-style constraints remaining rare. Third, RQ3 provides controlled evidence that large shifts in whether agents write tests translate into only small shifts in task outcomes for most tasks, while efficiency effects can be substantial—inducing tests can increase interaction overhead without improving resolution, whereas suppressing tests can materially reduce API calls and token usage with only modest success losses.
Contributions.
The contribution of this work can be summarized as:
@@
  • A feedback-signal analysis of agent-written tests with a four-category assertion categorization. We separate verification-oriented assertions from observational outputs and introduce a rule-based AST classifier that maps assertions into four assertion categories. We find that tests largely serve an observational role: value-revealing prints consistently outnumber assertions; assertion usage is dominated by local-property and exact-value checks, whereas relational/range-style constraints are uncommon.
  • A causal evaluation of agent-written tests on task resolution. Through controlled prompt interventions that either encourage or suppress writing new test files, we quantify the causal effects of agent-written tests on task success and interaction efficiency. We demonstrate that large flips in test-writing status translate into only small changes in resolution outcomes for most tasks, whereas efficiency effects can be substantial—inducing tests can increase token and interaction overhead without improving success, while suppressing tests yields large cost reductions with only modest success drops.
Paper organization.
Section 2 describes our study setting, data collection, and measurement procedures. Section 3 characterizes emergent agent-written testing behaviors as process events—whether agents write tests, when they appear, and how intensively they are executed (RQ1). Section 4 shifts to test content by analyzing what feedback signals agent-written tests produce at execution time (assertions vs. value-revealing prints) and what types of assertions they use (RQ2). Section 5 evaluates the outcome and efficiency impacts of agent-written tests via prompt interventions that induce or suppress test writing (RQ3). Section 6 discusses implications, limitations, and future directions. Section 7 reviews related work, and Section 8 concludes the paper.
@@ 2. Methodology
In this section, we introduce the methodology for this study, including the benchmark, studied agent and LLMs, the extraction of agent-written tests, and implementation details. Our study is guided by three research questions:
  • RQ1: What Testing Behaviors Emerge Under a Light Agent Scaffold?
  • RQ2: What Feedback Signals Do Agent-Written Tests Provide, and What Types of Assertions Do They Use?
  • RQ3: Do Agent-Written Tests Truly Affect Task Resolution?
@@ 2.1. Benchmark
We use SWE-bench Verified as a standardized source of real-world software-engineering trajectories in which agents autonomously write and use tests when solving a GitHub issue (Jimenez et al., 2024). SWE-bench Verified is a filtered subset of SWE-bench that contains 500 instances. Each benchmark instance provides a GitHub issue, a fixed repository snapshot, and the official evaluation harness. We analyze agent-written test artifacts within the observed trajectories (OpenAI, 2024).
2.2. Agent and its LLMs
While many recent LLM-based agents incorporate curated testing components, such as specialized validation modules, dedicated test-planning stages, or multi-agent coordination (Liu et al., 2024; Zhang et al., 2024; Ruan et al., 2024; Cognition Labs, 2024), these frameworks can confound a model’s intrinsic tendencies with scaffold-induced constraints. To better isolate base-model behavior, we adopt mini-SWE-agent (SWE-bench Team, 2024; SWE-agent Team, 2024). It provides a lightweight agent work loop restricted to a standard bash interface: the agent interacts with the repository solely through the bash tool, executing commands in a bash shell (e.g., running python) and using standard command-line utilities to inspect and modify files. The model can create executable Python test files on the fly and run them via bash as part of its workflow. Crucially, mini-SWE-agent does not provide additional testing-specific functions or dedicated testing tools (e.g., test planners or structured testing modules), so the testing decisions (whether, when, and how) are left to the model. Accordingly, any observed behaviors (e.g., creating or running test artifacts) can be interpreted as model-native ones.
We select a diverse set of strong LLMs to capture heterogeneous agent-written testing behaviors under mini-SWE-agent. Specifically, we reference the SWE-bench Bash Only leaderboard [^1] https://www.swebench.com/ as of the time of writing (2025-12-11) and identify the top-six model families (rather than top entries, since a family may appear with multiple variants on the leaderboard). For each family, we use its highest-ranked model as the representative: claude-opus-4.5 (Anthropic, 2025) (74.4%), gemini-3-pro-preview (Google Cloud, 2025) (74.2%), gpt-5.2 (OpenAI, 2025) (71.8%), kimi-k2-thinking (Moonshot AI, 2025) (63.4%), minimax-m2 (MiniMax, 2025) (61.0%), and deepseek-v3.2-reasoner (DeepSeek, 2025a) (60.0%).
@@ 2.3. Data Extraction of Agent-Written Tests
In our study, agent-written tests are test files that an agent writes using the bash tool during task resolution. We extract agent-written tests from task trajectories, which are time-ordered interaction logs recorded during issue resolution that include the agent’s intermediate reasoning, concrete actions (e.g., bash commands), and the resulting observations. To find test files written during a trajectory, we scan the logged bash actions for file-writing operations, most commonly here-doc writes such as cat <<’EOF’ > path/to/file.pyEOF. We then keep only files whose paths match common Python test naming patterns, including filenames that start with test_ or end with _test.py or tests.py.
@@ 2.4. Implementation Details
We run all experiments using the official mini-SWE-agent codebase. All tasks are executed on a Linux server (Ubuntu 22.04.5) with an AMD Ryzen Threadripper PRO 7975WX CPU (32 cores / 64 threads), 251 GiB RAM. For model inference, we access the four LLMs through a combination of official provider APIs and the OpenRouter API. For evaluation, we use the official SWE-bench sb-cli tool to score each submitted patch under the benchmark harness. Across all experiments reported in this paper, the total LLM API cost is approximately USD 1,600.
@@ 3. RQ1: What Testing Behaviors Emerge Under a Light Agent Scaffold?
Model #Tasks Tasks w/ tests (count, %) Mean #tests #Tasks Tasks w/ tests (count, %) Mean #tests #Tasks Tasks w/ tests (count, %) Mean #tests
claude-opus-4-5 372 314 (84.4%) 3.33 128 101 (78.9%) 4.12 500 415 (83.0%) 3.52
gemini-3-pro-preview 371 235 (63.3%) 2.02 129 73 (56.6%) 2.16 500 308 (61.6%) 2.05
@@
kimi-k2-thinking 317 309 (97.5%) 3.48 183 178 (97.3%) 3.83 500 487 (97.4%) 3.61
minimax-m2 305 302 (99.0%) 4.82 195 191 (97.9%) 5.76 500 493 (98.6%) 5.19
deepseek-v3.2-reasoner 300 277 (92.3%) 3.55 200 169 (84.5%) 4.08 500 446 (89.2%) 3.75
Notes. “Tasks w/ tests (count, %)” reports the number (and percentage) of tasks that write at least one test artifact within each outcome split. “Mean #tests” reports the mean number of distinct test artifacts, computed only over tasks that write at least one test artifact.
@@ Results.
Table 1 shows that writing tests is common for most models, but not for gpt-5.2. Some models write tests in almost every task (e.g., minimax-m2 and kimi-k2-thinking). In contrast, gpt-5.2 almost never writes tests (3/500 tasks). Within the same model, resolved and unresolved tasks usually have similar test-writing rates. When tests are written, unresolved tasks often write as many or more distinct test artifacts than resolved tasks. This may reflect that harder tasks trigger more trial-and-error.
RQ1.1 Test Writing: Key Pattern Test writing is widespread across models in our high-autonomy issue-resolution setting. Most models write tests in a majority of tasks (61.6–98.6%), and four of six do so in at least 83% of tasks; gpt-5.2 is a clear outlier with near-zero test writing (0.6%).
@@
Model #Tasks w/ tests First test-writing position Last test-writing position Test-writing span #Tasks w/ tests First test-writing position Last test-writing position Test-writing span
claude-opus-4-5 314 0.34 0.75 0.41 101 0.30 0.78 0.48
gemini-3-pro-preview 235 0.53 0.67 0.14 73 0.55 0.70 0.15
kimi-k2-thinking 309 0.40 0.82 0.42 178 0.40 0.82 0.42
minimax-m2 302 0.35 0.86 0.51 191 0.29 0.85 0.56
deepseek-v3.2-reasoner 277 0.43 0.80 0.37 169 0.40 0.80 0.40
All models 1440 0.40 0.78 0.38 712 0.37 0.80 0.43
Notes. Values are macro-over-tasks means. “First test-writing position” and “Last test-writing position” are the normalized positions of the first and last test-writing events relative to total_steps. “Test-writing span” is the normalized distance between the first and last test-writing positions.
@@
Table 2 summarizes test-writing positions for tasks that write tests. Across all models, the average first test-writing position is 0.40 for resolved tasks and 0.37 for unresolved tasks. The average last test-writing position is 0.78 (resolved) and 0.80 (unresolved). Models differ in when they start writing tests. For example, gemini-3-pro-preview starts later (0.53–0.55), while minimax-m2 and claude-opus-4-5 start earlier (0.29–0.35). Most models finish test writing late in the task (last position around 0.75–0.86). Models also differ in how spread out test writing is. gemini-3-pro-preview has a short span (0.14–0.15). minimax-m2 has a wider span (0.51–0.56), and claude-opus-4-5 is also relatively wide (0.41–0.48). kimi-k2-thinking is almost identical between resolved and unresolved tasks (0.40–0.82; span 0.42). Overall, unresolved tasks have a slightly larger average span than resolved tasks (0.43 vs. 0.38).
RQ1.2 Test-Writing Timing: Key Pattern Test writing typically finishes late, but its start time and span are mainly model-dependent; unresolved tasks are only slightly more spread out (0.43 vs. 0.38).
3.3. RQ1.3 Execution: How Intensively Are Agent-Written Tests Executed, and With What Process Outcomes?
@@
Model #Tasks w/ tests Mean ExecCount Mean ExecPerTest Mean FailRate (%) #Tasks w/ tests Mean ExecCount Mean ExecPerTest Mean FailRate (%)
claude-opus-4-5 314 4.87 1.50 11.97 101 6.27 1.68 11.14
gemini-3-pro-preview 235 2.71 1.51 8.53 73 2.79 1.40 7.08
kimi-k2-thinking 309 5.39 1.62 24.95 178 6.54 1.76 21.05
minimax-m2 302 7.19 1.55 24.11 191 9.70 2.09 24.10
deepseek-v3.2-reasoner 277 3.74 1.11 27.37 169 4.66 1.32 29.55
All models 1440 4.89 1.46 19.68 712 6.52 1.70 21.05
Notes. #Tasks: tasks with test writing. Mean ExecCount: test executions per task. Mean ExecPerTest: executions per agent-written test artifact. Mean FailRate: % executions with non-zero return codes (macro-over-tasks).
@@
Table 3 summarizes execution effort and process-level outcomes for tasks that write tests. Across all models, unresolved tasks execute tests more often than resolved tasks (Mean ExecCount: 6.52 vs. 4.89). They also rerun tests more per written test artifact (Mean ExecPerTest: 1.70 vs. 1.46). FailRate is slightly higher for unresolved tasks (21.05% vs. 19.68%). Models differ strongly in execution intensity. gemini-3-pro-preview runs tests the least (ExecCount $\approx$ 2.7–2.8). minimax-m2 runs tests the most, especially for unresolved tasks (ExecCount 9.70; ExecPerTest 2.09). FailRate also varies by model. claude-opus-4-5 and gemini-3-pro-preview have lower FailRate (about 7–12%), while deepseek-v3.2-reasoner, kimi-k2-thinking, and minimax-m2 are higher (about 21–30%).
RQ1.3 Test Execution: Key Pattern For tasks that write tests, unresolved tasks run tests more often and rerun them more per written test artifact (ExecCount 6.52 vs. 4.89; ExecPerTest 1.70 vs. 1.46), while process-level execution failures vary mainly by model (FailRate $\sim$7%–30%).
@@
RQ1 provides a descriptive baseline of agent-written testing behaviors in a high-autonomy setting. Test writing is common for most models, but extremely rare for gpt-5.2 (Table 1). For tasks that write tests, test writing usually ends late in the task, while the start position and the writing span vary mainly by model (Table 2). Unresolved tasks are slightly more spread out in test writing on average (span 0.43 vs. 0.38). Finally, among tasks that write tests, unresolved tasks execute tests more often and rerun them more per written test artifact (Table 3). Process-level execution failures also vary by model (about 7–30% FailRate). RQ1, however, only describes when and how often tests appear and run, not what feedback they provide. RQ2 therefore analyzes the feedback encoded in agent-written tests.
4. RQ2: What Feedback Signals Do Agent-Written Tests Provide, and What Types of Assertions Do They Use?

### v2

@@ Abstract.
Large Language Model (LLM) code agents increasingly resolve repository-level issues by iteratively editing code, invoking tools, and validating candidate patches. In these workflows, agents often write tests on the fly, but the value of this behavior remains unclear. For example, GPT-5.2 writes almost no new tests yet achieves performance comparable to top-ranking agents. This raises a central question: do such tests meaningfully improve issue resolution, or do they mainly mimic a familiar software-development practice while consuming interaction budget?
To better understand the role of agent-written tests, we analyze trajectories produced by six strong LLMs on SWE-bench Verified. Our results show that test writing is common, but resolved and unresolved tasks within the same model exhibit similar test-writing frequencies. When tests are written, they mainly serve as observational feedback channels, with value-revealing print statements appearing much more often than assertion-based checks. Based on these insights, we perform a prompt-intervention study by revising the prompts used with four models to either increase or reduce test writing. The results suggest that prompt-induced changes in the volume of agent-written tests do not significantly change final outcomes in this setting. Taken together, these results suggest that current agent-written testing practices reshape process and cost more than final task outcomes.
@@
journalyear: 2026conference: ; October, 2026; Washington, DC, USA
@@ 1. Introduction
Code agents increasingly tackle software issues by combining Large Language Models (LLMs) (OpenAI, 2025; DeepSeek, 2025b; Chen and Jiang, 2024) with tools and interaction protocols that let them edit real repositories, invoke tools, and attempt end-to-end issue resolution (Yang et al., 2024a; Örwall, 2024; Wang et al., 2024b; Gao et al., 2025b; Hong et al., 2024; Qian et al., 2024; Ehrlich et al., 2025; Xia et al., 2024). In this paper, a code agent denotes an LLM coupled with external tools and an iterative action–observation loop, and the scaffold refers to the surrounding tool interface and interaction protocol that specifies the agent’s allowed actions and feedback. Among the diverse skills required by code agents, testing plays a critical role: it exposes regressions, validates hypotheses, and provides a feedback loop during patch development (Zhang et al., 2024; Xia et al., 2024; Liu et al., 2024; Mu et al., 2025).
Figure 1. Overview of the study design. RQ1 examines testing behaviors, RQ2 analyzes feedback signals in agent-written tests, and RQ3 studies observed outcome and efficiency changes under test-writing interventions.
@@
To better understand agent-written tests, we conduct a quantitative analysis of agent trajectories on SWE-bench Verified (OpenAI, 2024) using mini-SWE-agent (SWE-agent Team, 2024), where testing is optional and not enforced by any hard-coded procedure. Agent-written testing is prevalent for several strong models. For example, Claude Opus 4.5 (ranked #1 in this setting, 74.4% resolution) generates at least one new test artifact in about 83% of tasks. By contrast, GPT-5.2 achieves a comparable resolution rate (71.8%), only 2.6 percentage points below Claude Opus 4.5, while generating near-zero new tests (only in 0.6% of tasks). This observation motivates a core question: when prompts and model tendencies lead agents to write more tests, do the observed outcomes actually improve, or do models merely mimic a learned software-development practice while the resulting tests contribute little to the final patch? If the latter holds, then the widespread creation and execution of agent-written tests may represent a substantial waste of resources, consuming interaction budget without meaningful gains in task success. Therefore, we argue that a systematic empirical study is needed to understand the role of agent-written tests in resolving software issues.
Prior work mostly evaluates and benchmarks LLM-generated tests under predefined testing objectives and fixed quality metrics (e.g., unit tests, assertions, or issue-reproducing tests), typically with respect to a fixed target program or snapshot of code under test (Lops et al., 2025; Zhang et al., 2025b; Mündler et al., 2024; Yuan et al., 2024; Wang et al., 2025, 2024a; Schäfer et al., 2023). However, in complex real-world GitHub issue resolution (Jimenez et al., 2024; Khandpur et al., 2025), the codebase and candidate patches evolve over time, and test writing and usage arise dynamically as self-directed behaviors rather than pre-specified evaluation objectives. Yet the intrinsic tendency of high-autonomy agents to write and use tests during such issue resolution—and how prompt-induced changes in test writing relate to observed resolution outcomes and costs—has not been systematically studied. This motivates a closer empirical investigation of agent-written tests, guided by the following three research questions.
@@
Guided by this gap, we study the role of agent-written tests in GitHub issue resolution. Figure 1 summarizes our study design, which decomposes the problem into three complementary research questions. RQ1 characterizes the agents’ testing behaviors under a light scaffold where test writing is optional: whether agents write tests, when they introduce them, and how intensively they execute them. RQ2 shifts from behaviors to test content, investigating what feedback signals agent-written tests actually emit at execution time (assertions vs. value-revealing prints), what types of assertions they use, and what kinds of runtime information those value-revealing prints typically inspect. RQ3 examines observed outcome and efficiency changes: by revising prompts to encourage or discourage writing new tests, we measure how changing test-writing behavior is accompanied by changes in task resolution outcomes and efficiency costs (API calls and tokens).
@@
Across models, agent-written testing is best understood as a model-dependent process style rather than a dependable driver of success. RQ1 shows that test writing is widespread but only weakly aligned with success. RQ2 shows that agent-written tests primarily serve as an observational feedback channel, with value-revealing prints dominating assert-based checks. RQ3 shows that prompt-induced changes in test writing have only small observed effects on task outcomes for most tasks, but can materially change efficiency.
The contributions of this work can be summarized as follows:
@@
  • A feedback-signal analysis of agent-written tests covering assertions and value-revealing prints. We separate verification-oriented assertions from observational outputs and introduce rule-based AST analyses that map assertions into four categories and value-revealing prints into three coarse categories. We find that tests largely serve an observational role: value-revealing prints consistently outnumber assertions, those prints are dominated by value/content inspection and exception/status signals, and assertion usage is dominated by local-property and exact-value checks.
  • A prompt-intervention study of agent-written tests on observed outcomes and efficiency. Through controlled prompt interventions that either encourage or suppress writing new test files, we study how changes in test-writing cues relate to observed task success and interaction efficiency under the same agent scaffold. We show that large flips in test-writing status are accompanied by only small observed changes in resolution outcomes for most tasks, whereas efficiency changes can be substantial. Inducing tests can increase token and interaction overhead without improving success, while suppressing tests yields large cost reductions with only modest success drops.
@@ 2. Methodology
In this section, we introduce the methodology for this study, including the benchmark, the studied agent and LLMs, the extraction of agent-written tests, and implementation details. Our study is guided by three research questions: RQ1: What Testing Behaviors Emerge Under a Light Agent Scaffold? RQ2: What Feedback Signals Do Agent-Written Tests Provide? RQ3: How Does Prompting Test Writing Change Observed Outcomes and Costs?
@@ 2.1. Benchmark
We use SWE-bench Verified as our benchmark. The original SWE-bench benchmark is built from resolved GitHub issues drawn from 12 open-source Python repositories (Jimenez et al., 2024). SWE-bench Verified is a 500-instance subset released after a human-screening effort led by OpenAI in collaboration with the SWE-bench authors (OpenAI, 2024). Each instance provides a GitHub issue, a fixed repository snapshot, and the official evaluation harness. We analyze agent-written test artifacts within the resulting trajectories.
$2.$ Agent and its LLMs
While many recent LLM-based agents incorporate curated testing components, such as specialized validation modules, dedicated test-planning stages, or multi-agent coordination (Liu et al., 2024; Zhang et al., 2024; Ruan et al., 2024; Cognition Labs, 2024), these frameworks can confound a model’s intrinsic tendencies with scaffold-induced constraints. To better isolate base-model behavior, we adopt mini-SWE-agent (SWE-bench Team, 2024; SWE-agent Team, 2024). It provides a lightweight agent work loop restricted to a standard bash interface: the agent interacts with the repository solely through the bash tool, executing commands in a bash shell (e.g., running python) and using standard command-line utilities to inspect and modify files. The model can create executable Python test files on the fly and run them via bash as part of its workflow. Crucially, although the default mini-SWE-agent prompt includes a brief natural-language recommendation such as “Test edge cases to ensure your fix is robust,” this instruction is only advisory: it does not introduce any testing-specific function, dedicated testing tool, or hard-coded workflow component (e.g., a test planner, structured testing module, or enforced test-execution stage). Thus, in our setting, testing remains optional at runtime: the model may follow, delay, or ignore that recommendation, and the decisions of whether, when, and how to test are still left to the model. Accordingly, any observed behaviors (e.g., creating or running test artifacts) can be interpreted as model-native ones.
We select a diverse set of strong LLMs to capture heterogeneous agent-written testing behaviors under mini-SWE-agent. For model selection, we use the SWE-bench Bash Only leaderboard with a cutoff date of 2025-12-11111https://www.swebench.com/ and identify the top-six model families (rather than top entries, since a family may appear with multiple variants on the leaderboard). For each family, we use its highest-ranked model as the representative: claude-opus-4.5 (Anthropic, 2025) (74.4%), gemini-3-pro-preview (Google Cloud, 2025) (74.2%), gpt-5.2 (OpenAI, 2025) (71.8%), kimi-k2-thinking (Moonshot AI, 2025) (63.4%), minimax-m2 (MiniMax, 2025) (61.0%), and deepseek-v3.2-reasoner (DeepSeek, 2025a) (60.0%). In the remainder of the paper, we refer to these models as Claude Opus 4.5, Gemini 3 Pro, GPT-5.2, Kimi K2 Thinking, MiniMax M2, and DeepSeek v3.2 Reasoner. In tables and figures, we further shorten Claude Opus 4.5, Kimi K2 Thinking, and DeepSeek v3.2 Reasoner to Claude 4.5, Kimi K2-T, and DeepSeek v3.2-R to save space. In each case, the shortened form still uniquely identifies the official model version evaluated in this study, without introducing ambiguity with another model release.
@@ 2.3. Data Extraction of Agent-Written Tests
In our study, agent-written tests are test files that an agent writes using the bash tool during task resolution. We extract agent-written tests from task trajectories, which are time-ordered interaction logs recorded during issue resolution that include the agent’s intermediate reasoning, concrete actions (e.g., bash commands), and the resulting observations. To find test files written during a trajectory, we scan the logged bash actions for file-writing operations, most commonly here-doc writes such as cat <<’EOF’ > path/to/file.pyEOF. We then keep only files whose paths match common Python test naming patterns, including filenames that start with test_ or end with _test.py or tests.py (pytest developers, 2026).
@@ 2.4. Implementation Details
We run all experiments using the official mini-SWE-agent codebase. All tasks are executed on a Linux server (Ubuntu 22.04.5) with an AMD Ryzen Threadripper PRO 7975WX CPU (32 cores / 64 threads), 251 GiB RAM. For model inference, we access LLMs through a combination of official provider APIs and the OpenRouter API. For evaluation, we use the official SWE-bench sb-cli tool to score each submitted patch under the benchmark harness. Across all experiments reported in this paper, the total LLM API cost is approximately USD 1,600.
@@ 3. RQ1: What Testing Behaviors Emerge Under a Light Agent Scaffold?
Model #Tasks Tasks w/ tests Mean #tests #Tasks Tasks w/ tests Mean #tests #Tasks Tasks w/ tests Mean #tests
Claude 4.5 372 314 (84.4%) 3.33 128 101 (78.9%) 4.12 500 415 (83.0%) 3.52
Gemini 3 Pro 371 235 (63.3%) 2.02 129 73 (56.6%) 2.16 500 308 (61.6%) 2.05
@@
Kimi K2-T 317 309 (97.5%) 3.48 183 178 (97.3%) 3.83 500 487 (97.4%) 3.61
MiniMax M2 305 302 (99.0%) 4.82 195 191 (97.9%) 5.76 500 493 (98.6%) 5.19
DeepSeek v3.2-R 300 277 (92.3%) 3.55 200 169 (84.5%) 4.08 500 446 (89.2%) 3.75
Notes. “Tasks w/ tests” reports count and percentage within each outcome split. “Mean #tests” is computed only over tasks that write at least one test artifact.
@@ Results.
Table 1 shows that writing tests is common for most models, but not for GPT-5.2. Some models write tests in almost every task (e.g., MiniMax M2 and Kimi K2 Thinking). In contrast, GPT-5.2 almost never writes tests (3/500 tasks). Within the same model, resolved and unresolved tasks usually have similar test-writing rates. When tests are written, unresolved tasks often write as many or more distinct test artifacts than resolved tasks. This may reflect that harder tasks trigger more trial-and-error.
RQ1.1 Test Writing: Key Pattern Test writing is common for most models in this high-autonomy setting, but GPT-5.2 is a clear outlier with near-zero test writing.
@@
Model #Tasks First pos. Last pos. Span #Tasks First pos. Last pos. Span
Claude 4.5 314 0.34 0.75 0.41 101 0.30∗∗∗ 0.78∗∗ 0.48∗∗
Gemini 3 Pro 235 0.53 0.67 0.14 73 0.55∗∗∗ 0.70∗∗∗ 0.15∗∗∗
Kimi K2-T 309 0.40 0.82 0.42 178 0.40∗∗∗ 0.82∗∗∗ 0.42∗∗∗
MiniMax M2 302 0.35 0.86 0.51 191 0.29∗∗∗ 0.85∗∗ 0.56∗∗
DeepSeek v3.2-R 277 0.43 0.80 0.37 169 0.40∗∗∗ 0.80∗∗∗ 0.40∗∗∗
All models 1440 0.40 0.78 0.38 712 0.37∗∗∗ 0.80∗∗∗ 0.43∗∗∗
Notes. Macro-over-tasks means over tasks with tests. Asterisks mark significant resolved-vs.-unresolved differences within a model (two-sided Mann–Whitney U: $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$).
@@
Table 2 summarizes test-writing positions for tasks that write tests. Across all models, the average first test-writing position is 0.40 for resolved tasks and 0.37 for unresolved tasks. The average last test-writing position is 0.78 (resolved) and 0.80 (unresolved). Models differ in when they start writing tests. For example, Gemini 3 Pro starts later (0.53–0.55), while MiniMax M2 and Claude Opus 4.5 start earlier (0.29–0.35). Most models finish test writing late in the task (last position around 0.75–0.86). Models also differ in how spread out test writing is. Gemini 3 Pro has a short span (0.14–0.15). MiniMax M2 has a wider span (0.51–0.56), and Claude Opus 4.5 is also relatively wide (0.41–0.48). Kimi K2 Thinking is almost identical between resolved and unresolved tasks (0.40–0.82; span 0.42). Overall, unresolved tasks have a slightly larger average span than resolved tasks (0.43 vs. 0.38).
RQ1.2 Test-Writing Timing: Key Pattern Test writing typically finishes late, while its start time and span are mainly model-dependent; unresolved tasks are only slightly more spread out.
$3.$ RQ1.3 Execution: How Intensively Are Agent-Written Tests Executed, and With What Process Outcomes?
@@
Model #Tasks w/ tests Exec Count Exec PerTest FailRate (%) #Tasks w/ tests Exec Count Exec PerTest FailRate (%)
Claude 4.5 314 4.87 1.50 11.97 101 6.27∗∗∗ 1.68 11.14
Gemini 3 Pro 235 2.71 1.51 8.53 73 2.79 1.40 7.08
Kimi K2-T 309 5.39 1.62 24.95 178 6.54 1.76 21.05
MiniMax M2 302 7.19 1.55 24.11 191 9.70∗∗∗ 2.09∗∗ 24.10
DeepSeek v3.2-R 277 3.74 1.11 27.37 169 4.66 1.32 29.55
All models 1440 4.89 1.46 19.68 712 6.52∗∗∗ 1.70∗∗ 21.05
Notes. #Tasks: tasks with test writing; all other values are macro-over-tasks means. Asterisks mark significant resolved-vs.-unresolved differences within a model (two-sided Mann–Whitney U: $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$).
@@
Table 3 summarizes execution effort and process-level outcomes for tasks that write tests. Across all models, unresolved tasks execute tests more often than resolved tasks (Mean ExecCount: 6.52 vs. 4.89). They also rerun tests more per written test artifact (Mean ExecPerTest: 1.70 vs. 1.46). Mann–Whitney U tests show that these aggregate resolved-vs.-unresolved differences are statistically significant for ExecCount and ExecPerTest, but not for FailRate. At the model level, significant differences appear mainly for Claude Opus 4.5 (higher ExecCount in unresolved tasks) and MiniMax M2 (higher ExecCount and ExecPerTest in unresolved tasks). FailRate is slightly higher for unresolved tasks (21.05% vs. 19.68%), but this aggregate difference is not statistically significant. Models differ strongly in execution intensity. Gemini 3 Pro runs tests the least (ExecCount $\approx$ 2.7–2.8). MiniMax M2 runs tests the most, especially for unresolved tasks (ExecCount 9.70; ExecPerTest 2.09). FailRate also varies by model. Claude Opus 4.5 and Gemini 3 Pro have lower FailRate (about 7–12%), while DeepSeek v3.2 Reasoner, Kimi K2 Thinking, and MiniMax M2 are higher (about 21–30%).
RQ1.3 Test Execution: Key Pattern For tasks that write tests, unresolved runs execute them more intensively, and these aggregate differences are statistically significant, while process-level execution failures vary mainly by model.
@@
RQ1 shows that agent-written testing is better understood as a model-dependent process behavior than as a simple marker of eventual success. This, in turn, raises a more informative question for RQ2: when these tests do appear, what kind of feedback do they actually provide?
4. RQ2: What Feedback Signals Do Agent-Written Tests Provide?